A farmer divides 240 cattle among his 3 children in the ratio 5:4:3 how many cattle does each get?
2).if five men dig a drain in 14days how long would seven men take?

Answer by Boreal(15235)

5:4:3 is the ratio
a multiple is 5x:4x:3x
That sum, 12x, equals 240
x=20
so it is 100, 80, 60
-
5 men can dig 1/14 of a drain a day or each man can dig 1/70 of a drain a day
7 men would then dig 7/70 or 1/10 of a drain a day.
So it would take 10 men to do it.
time is inversely proportional to number of men
t=k/men
14=k/5
k=70
t=k/7=70/7=10 days